# Colorful Rings of Partitions

### v1

@@ Abstract.
We visualize the identity $p(n)=\sum_{k=1}^{n}\sigma(k)p(n-k)/n$ for the partition function $p(n)$ involving the divisor function $\sigma$, add comments on the history of visualizations of numbers, illustrate how different mathematical fields play together when proving $\lim_{n\to\infty}p(n)^{1/n}=1$ and introduce finite or infinite rings of partitions.
@@ 1. Introduction
As far as we know, the earliest mathematicians worked with the help of manipulatives in the form of sticks, pebbles and ropes. Finger or toe cardinalities might have initiated the decimal or base 20 systems, smoothness the hexadecimal number system. Marks were carved into bones, clay, stone or bark [29, 19, 5]. Numbers were then written on Clay tablets or Papyrus or encoded in knots [45]. While historical discoveries are correlated to pedagogical time lines [14], history is always an effective ingredient for teaching [30]. Manipulatives are not only relevant historically or for teaching, they help to build intuition, even in a time of smartphones, tablets, virtual reality and artificial intelligence. The use of physical manipulative has flared up in the 3D printing age [43, 56]. Dubbed as the forth industrial evolution [52], it was recently over shadowed by the AI revolution [64]. Manipulatives will reappear in virtual reality set-ups, once interfaces become fully VARK: visual, auditory, read related or kinesthetic [15].
Figure 1. Partitions like the $p(4)=5$ partitions of $5$ generalize numbers. There is addition and multiplication in the ring $(\mathcal{P},+,*,0,1)$. Note that $1*1=1$ but $1+1=(1,1)\neq(2)$. While $(2,2)=(2)+(2)$, the partition $(4)$ is not the sum of smaller partitions; it is an additive partition prime. Every $n=(n_{1},\dots,n_{k})$ with either $|n|=\sum_{j=1}^{k}n_{j}$ or $k$ a rational prime is a multiplicative partition prime but there are more: $(1,2,3,4)$ is a multiplicative partition prime even so $|n|=1+2+3+4$, and $k=4$ are not prime. Modular partition arithmetic exists in finite rings $\mathcal{P}_{n}$ in which numbers are of the form $(x_{1},\dots,x_{k}),x_{j}\in\{1,\dots,n-1\},x_{i}\leq x_{i+1}$, where $(x_{1}=1,\dots,x_{n}=1)=0$ and $(n_{1},\dots,n_{k-1},n)=(n_{1},\dots,n_{k-1})$. As on elliptic curves we get Abelian groups allowing cryptology.
@@ 2. Partitions as numbers
To illustrate how manipulatives can lead to innovation we look at partitions as ”generalized numbers” which are more fundamental than “natural numbers” that initiate the building $\mathbb{N}\subset\mathbb{Z}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\subset\mathbb{H}\subset\mathbb{O}$. We can add finite simple graphs using the join operation [65]. After completing the monoid $(G,\oplus)$ to a group, it leads with the large multiplication [53] to the Sabidussi ring $(\mathbb{G},\oplus,\otimes,0,1)$ in which the empty graph is the zero element and the $1$-point graph $1$ is the unit. This ring is isomorphic to the Shannon ring $(\mathcal{Z},+,*,0,1)$ in which $+$ is the disjoint union and * is the Shannon multiplication [57]. See [21, 58, 60, 36, 37, 38].
@@
The ring of partitions $\mathcal{P}$ is a subring of the Sabidussi ring $\mathcal{S}$ and are represented by multi-partite graphs, elements generated by zero dimensional subgraphs. The integers $\mathcal{Z}$ are the subring of the Sabidussi ring generated by complete graphs. As for the partition ring: the concept of a number like ”5” is now much richer. It contains $7$ flavors of $5$ which are the $7$ partitions of $5$. The addition is the concatenation. This comes up naturally in the Cuisenaire picture [9]. The multiplication can be visualized geometrically also.
@@
The ring of partitions $\mathcal{P}$ is an extension of the ring $\mathbb{Z}$ of integers and contained in the Sabidussi ring of graphs $\mathcal{G}$ which is isomorphic to the Shannon ring of graphs $\mathcal{Z}$. A partition $n$ that adds up to a rational prime $|n|$ or contains a rational prime $k(n)$ number of elements is a multiplicative partition prime. For every rational $n$, there is a modular partition ring $\mathcal{P}_{n}$. The number of elements in $\mathcal{P}_{n}$ grows exponentially in $n$. The maps $n\to|n|,n\to k(n)$ are ring homormorphisms from $\mathcal{P}$ to $\mathbb{Z}$.
@@
$\mathcal{Z}$ is generated by complete graphs. $\mathcal{P}$ is generated by zero-dimensional graphs. Note that $K_{1}\oplus K_{1}\cdots\oplus K_{1}=K_{n}$ is a complete graph so that $\mathcal{Z}\subset\mathcal{P}$. $\mathcal{G}$ is generated by all graphs. The duality in the form of graph complement represents $\mathcal{Z}$ in the Shannon ring as generated by zero dimensional graphs and $\mathcal{P}$ generated by complete graphs and $\mathcal{G}$ generated by all graphs. To see the prime statement, note that both $n=(n_{1},\dots,n_{k})\to|n|=\sum_{j=1}^{k(p)}n_{j}$ and $n\to k(n)$ are ring homomorphisms from $\mathcal{P}$ to $\mathbb{Z}$. To see that there is a modular arithmetic $\mathcal{P}_{n}$, we not only just assume that the parts $n_{k}$ of the partition are in $\{1,\dots,n-1\}$ by subtracting $n$ if there should be $\geq n$ elements in a slot of the partition (removing the slot if it becomes zero), or by subtracing $1$ from each of first $n$ slots. This has the effect that there will be less than $n$ slots all the time. The procedure is uniquely defined as we have always $10=leqk\leq n-$ 1 for a partition $(n_{1},n_{2}\dots,n_{k})\in\mathcal{P}_{n}$. The partition $(2,1,3,4,1)$ would reduce to $(1,2,3)$ in $\mathcal{P}_{5}$. The homormorphism statements follow from $|n+m|=|n|+|m|,k(n+m)=k(n)+k(m),|n*m|=|n||m|,k(n*m)=|n||m|$. ∎
@@ 3. Additive number theory
We illustrate the power of colored sticks, we use a recursion theorem which was obtained in [31]. It was found independently with the help of Cuisenaire material. It turns out that the mathematical theorem result has been known since more than 100 years [49, 1]. It probably was known to Euler already. Let $\sigma(n)$ denote the divisor $\sigma$-function which gives the sum of the proper divisors of an integer $n$. The partition number $p(n)$ gives the number of different ways to write $n$ as a sum of smaller positive integers where the order of the summands does not matter.
@@
The arrangement of all partitions is a rectangle $P(n)$ of area $np(n)$. Let $f\downarrow P(n)$ denot how many times the stick $f$ appears in the pile. One can now take from every row containing an $f$ one piece away to see
@@
The theorem can be seen also as a fixed point theorem $p=T(p)$ for a convolution average
@@
The book title of the ATM resources [6] shows similar arrangements for partitions of numbers if the order matters. See [31] on page 133 in an other arrangement. [6] states ”After Gattegno Curriculum Graph 1974”. The idea of illustrating identities as such must have appeared first in [18] page 5, interlocking staircases illustrating $10=1+9=2+8+\dots=9+1=10$. We could not find any partition illustrations earlier than [31], also not in Gattegno’s original books [17], nor in [9].
@@
This is amazing because $\sigma(n)$ uses multiplicative properties of $n$ and and $p(n)$ uses additive structures. The initial conditions are different. Both assume $p(n)=\sigma(n)=0$ if $n<0$. The partition function starts with $(p(1),p(2))=(1,2)$ and takes $p(0)=1$. The divisor function starts with $(\sigma(1),\sigma(2))=(1,3)$ and uses $\sigma(n-n)=n$ in the recursion, when computing $p(n)$.
@@ 4. Dynamics
Hadamard’s lacunary theorem [26] about power series $\sum_{k}c_{k}z^{n_{k}}$ shows that if $n_{k}+1/n_{k}\geq r>1$, then $r=1$ is a natural bounday. This means that boundedness at one point $|z|=r$ would imply that the radius of convergence is larger than $r$. Define $g(\alpha)=(1-re^{2\pi i\alpha})^{-1}$ so that $f(\alpha)=\log(P(z))=-\sum_{k=1}^{\infty}g(k\alpha)$. For $r=|z|<1$, and $z=r\exp(2\pi i\alpha)$ with Diophantine $\alpha$, the theorem of Gottschalk-Hedlund [28] assures that $f$ is a coboundary, meaning $f(x)=F(x+\alpha)-F(x)$, implying that the sum is bounded. Having so proven that the Taylor series has no singularity on $|z|<1$, the partition function $p(n)$ satisfies $\limsup_{n}|p(n)|^{1/n}\geq 1$. The other inequality $\limsup_{n}|p(n)|^{1/n}\leq 1$ implies together with $p(n)\geq 1$ that $\sum_{n=1}^{\infty}p(n)z^{n}\geq\sum_{n=1}^{\infty}z^{n}$ so that the radius of convergence of the former is larger or equal than the radius of convergence of the later, which is $1$. ∎
@@ 5. Psychology
How do human learn, teach and do research? How do they gain intuition? Psychological aspects of teaching have been looked at in research [20] or education [48]. Hadamard stresses that visual thinking is key for intuition. Norton suggests that a number is defined through our own actions and mentions in the chapter ”What are numbers” cites historical definitions of numbers like Euclid’s notion ”A number is a multitude composed of units.” appearing in book VII. The development of numbers is also linked to language and notation [46]. A more topological approach to numbers are quipu, talking knots [45, 2].
@@
While commutative structures have been challenged since the emergence of quantum mechanics and under the umbrella of non-commutative mathematics [7], the commutativity axiom for the additive structure of the integers has never been put up to debate. But as argued in [40], there are reasons to consider a group generated by two involutions more fundamental than a group by a single generator $1$. The dihedral approach very much is in the philosophy of reflection geometry [3, 24] where reflections are used to generate the non-Abelian Euclidean group $\mathbb{R}^{n}\rtimes O(n)$ consisting of translations and orthogonal transformations. Reflections at affine planes are all involutions. dighedral counting enters geometry if we look at the group generated by the exterior derivative $d$ and its dual $d^{*}$. We have $d^{2}=0$ and $(d^{*})^{2}=0$.
@@
Also notation and implementation is important. Good notation is pays off in the long term by allowing to think more effectively. Leibniz’s notations and nomenclature in calculus for example was much persuasive than Newton’s notation. In the case of manipulatives, the analog of notation is how to use material, shape or color and labels implemented in texture. Katherine Stern used colored cubes around a similar time than Cuisenaire. But the colors did not carry intuitive meaning. Other manipulatives had writings on them, making them less elegant.
@@ 6. History
Figure 6. Hardy (played in the movie by Jeremy Irons) explaining the asymptotic formula for the partition function in the 2015 film: ”The man who knew infinity”.
@@ 7. Creativity
Partitions appear when defining discrete sub-manifolds of given manifolds [41]. While working on this emerged the picture that the addition of two partitions $n=(n_{1}\dots,n_{k})$, $m=(m_{1},\dots,m_{l})$, in the form of concatenation $n+m=(n_{1},\dots,n_{k},m_{1},\dots,m_{l})$ and the product by foiling out: $n*m=(n_{1}m_{1},n_{2}m_{1},\dots,n_{k}m_{1},\dots,n_{1}m_{k}\dots,n_{k}m_{l})$. For example $(3,4)+(1,1,2)=(3,4,1,1,2)$ and $(3,4)*(1,1,2)=(3,3,6,4,4,8)$. leads to the idea to associate to a partition $n=(n_{1},\dots,n_{k})$ the $k$-partite graph $K_{n_{1},\dots,n_{k}}$. The partition $5=K_{1,1,1,1,1}$ for example is $K_{5}$ and $6==K_{3,3}$ is the utility graph. The addition is the Zykov join and the multiplication is the Sabidussi multiplication. This ring structure on multi-partite graph is dual to the Shannon ring structure, where addition is disjoint union and multiplication is Shannon multiplication.
@@
The freedom which these teachers had, is rather unthinkable today. It has become the privilege of private institutions. The times of the “Black Mountain college” [55, 11], where rival methodologies were explored and experimental forms were practiced, are long over. (The college is featured in the 2013 novel ”The longest ride” by Nickolas Sparks, and also made it into a movie in 2015, where the free teaching style at the college is very well visible.)
@@
By the way, the mathematician Max Dehn, who was a student of Hilbert, had taught from 1945-1952 at the college and is buried there [59]. The Wikipedia article about him (which follows [59]) states He enjoyed the forested mountains found in Black Mountain, and would often hold class in the woods, giving lectures during hikes. His lectures frequently drifted off topic on tangents about philosophy, the arts, and nature and their connection to mathematics. He and his wife took part in community meetings and often ate in the dining room. They also regularly had long breakfasts with Buckminster Fuller and his wife. As as source, [50] is given.
@@
The following quotation given in [59] about math education is remarkable: All instruction, including mathematics, should be an end in itself, not, in the first place, means. Therefore in mathematics one should be so simple and elementary as possible so that the structure of mathematical thinking becomes clear. If this is achieved then the student will have less difficulty to see mathematical structures in physical phenomena or even in phenomena of biology and social life.
Figure 7. This photo credited to the photography critic Nancy Newhall (1908-1974) is titled ”Max Dehn at the Black Mountain College” is dated around 1947. The gelatin silver negative had been gifted by Christi Newhall to the Eastman Museum, where one can find a high resolution version. Nancy Newhall visited together with her husband Beaumont Newhall the Black Mountain College for three summers in the years 1946-1949. The Newhalls photographed during that time the campus and its people.
@@
As one can read The administrative cages have been rolled out more and more in higher education. We could still develop my calculus courses independently without oversight 10 years ago. Today, we live in a time, with much more constraints. Yes, it is safer like that, but lack of initiative and originality will get us replaced by bots much faster than anticipated.

### v2

# The Colorful Ring of Partitions

@@ Abstract.
We visualize the identity $p(n)=\sum_{k=1}^{n}\sigma(k)p(n-k)/n$ for the partition function $p(n)$ involving the divisor function $\sigma$, add comments on the history of visualizations of numbers, illustrate how different mathematical fields play together when proving $\lim_{n\to\infty}p(n)^{1/n}=1$ and introduce the modular ring of partitions.
@@ 1. Introduction
As far as we know, the earliest mathematicians worked with the help of manipulatives in the form of sticks, pebbles and ropes. Finger or toe cardinalities might have initiated the decimal or base 20 systems, divisibility considerations the hexadecimal number system. Marks were carved into bones, clay, stone or bark [29, 19, 5]. Numbers were then written on Clay tablets or Papyrus or encoded in knots [45]. While historical discoveries are correlated to pedagogical time lines [14], history is always an effective ingredient for teaching [30]. Manipulatives are not only relevant historically or for teaching, they help to build intuition, even in a time of smartphones, tablets, virtual reality and artificial intelligence. The use of physical manipulative has flared up in the 3D printing age [43, 56]. Dubbed as the forth industrial evolution [52], it was recently over shadowed by the AI revolution [64]. Manipulatives might reappear more in virtual reality set-ups, once interfaces become fully VARK: visual, auditory, read related or kinesthetic [15].
Figure 1. Partitions like the $p(4)=5$ partitions of $5$ generalize numbers. There is addition and multiplication in the ring $(\mathcal{P},+,*,0,1)$. Note that $1*1=1$ but $1+1=(1,1)\neq(2)$. While $(2,2)=(2)+(2)$, the partition $(4)$ is not the sum of smaller partitions; it is an additive partition prime. Every $n=(n_{1},\dots,n_{k})$ with either $|n|=\sum_{j=1}^{k}n_{j}$ or $k$ a rational prime is a multiplicative partition prime but there are more: $(1,2,3,4)$ is a multiplicative partition prime, even so $|n|=1+2+3+4$, and $k=4$ are not prime.
@@ 2. Partitions as numbers
To illustrate how manipulatives can lead to innovation, we look at partitions as “generalized numbers” enlarging “natural numbers” that initiate the build-up $\mathbb{N}\subset\mathbb{Z}\subset\mathbb{Q}\subset\mathbb{R}\subset\mathbb{C}\subset\mathbb{H}\subset\mathbb{O}$. The generalization is geometric: we can “add ”finite simple graphs using the join operation [65]. After completing the monoid $(\mathcal{G},\oplus)$ to a group, it leads with the large multiplication [53] to the Sabidussi ring $(\mathbb{G},\oplus,\otimes,0,1)$ in which the empty graph is the zero element and the $1$-point graph $1$ is the unit. This ring is isomorphic to the Shannon ring $(\mathcal{Z},+,*,0,1)$, in which $+$ is the disjoint union and * is the Shannon multiplication [57]. See [21, 58, 60, 36, 37, 38].
@@
The ring of partitions $\mathcal{P}$ is a subring of the Sabidussi ring $\mathcal{S}$. Its elements are multi-partite graphs, elements generated by zero dimensional subgraphs. The integers $\mathcal{Z}$ are the subring of the Sabidussi ring generated by complete graphs. In the partition ring, the concept of a number like ”5” is richer. It contains $7$ flavors of $5$, the $7$ partitions of $5$. The addition is the concatenation. This comes up naturally in the Cuisenaire picture [9]. The multiplication can be visualized geometrically also.
@@
The ring of partitions $\mathcal{P}$ is an extension of the ring $\mathbb{Z}$ of integers and contained in the Sabidussi ring of graphs $\mathcal{G}$ which is isomorphic to the Shannon ring of graphs $\mathcal{Z}$. A partition $n$ that adds up to a rational prime $p=|n|$ or contains a rational prime $k(n)=p$ number of elements is a multiplicative partition prime.
@@
$\mathcal{Z}$ is generated by complete graphs. $\mathcal{P}$ is generated by 00-dimensional graphs. Note that $K_{1}\oplus K_{1}\cdots\oplus K_{1}=K_{n}$ is a complete graph so that $\mathcal{Z}\subset\mathcal{P}$. $\mathcal{G}$ is generated by all graphs. The duality in the form of graph complement represents $\mathcal{Z}$ in the Shannon ring as generated by zero dimensional graphs and $\mathcal{P}$ generated by complete graphs and $\mathcal{G}$ generated by all graphs. To see the prime statement, note that both $n=(n_{1},\dots,n_{k})\to|n|=\sum_{j=1}^{k(p)}n_{j}$ and $n\to k(n)$ are ring homomorphisms from $\mathcal{P}$ to $\mathbb{Z}$. The homomorphism statements follow from $|n+m|=|n|+|m|,k(n+m)=k(n)+k(m),|n*m|=|n||m|,k(n*m)=|n||m|$. ∎
@@ 3. Additive number theory
We illustrate the power of colored sticks, we use a recursion theorem which was obtained in [31]. It was found independently with the help of Cuisenaire material. It turns out that the mathematical theorem result was known since more than 100 years [49, 1] and probably was known to Euler already. Let $\sigma(n)$ denote the divisor $\sigma$-function which gives the sum of the proper divisors of an integer $n$. The partition number $p(n)$ gives the number of different ways to write $n$ as a sum of smaller positive integers where the order of the summands does not matter.
@@
The arrangement of all partitions is a rectangle $P(n)$ of area $np(n)$. Let $f\downarrow P(n)$ denote how many times the stick $f$ appears in the pile. One can now take from every row containing an $f$ one piece away to see
@@
The theorem can be seen also as a fixed point result $p=T(p)$ for a convolution average
@@
The book title of the ATM resources [6] shows similar arrangements for partitions of numbers if the order matters. See also [31] (page 133). [6] states ”After Gattegno Curriculum Graph 1974”. The idea of illustrating identities as such must have appeared first in [18] page 5, interlocking staircases illustrating $10=1+9=2+8+\dots=9+1=10$. We could not find any partition illustrations earlier than [31], also not in Gattegno’s original books [17], nor in [9].
@@
This is amazing because $\sigma(n)$ uses multiplicative properties of $n$ and $p(n)$ uses additive structures. The initial conditions are different. Both assume $p(n)=\sigma(n)=0$ if $n<0$. The partition function starts with $(p(1),p(2))=(1,2)$ and takes $p(0)=1$. The divisor function starts with $(\sigma(1),\sigma(2))=(1,3)$ and uses $\sigma(n-n)=n$ in the recursion, when computing $p(n)$.
@@ 4. Dynamics
Hadamard’s lacunary theorem [26] about power series $\sum_{k}c_{k}z^{n_{k}}$ shows that if $n_{k}+1/n_{k}\geq r>1$, then $r=1$ is a natural boundary. This means that boundedness at one point $|z|=r$ would imply that the radius of convergence is larger than $r$. Define $g(\alpha)=(1-re^{2\pi i\alpha})^{-1}$ so that $f(\alpha)=\log(P(z))=-\sum_{k=1}^{\infty}g(k\alpha)$. For $r=|z|<1$, and $z=r\exp(2\pi i\alpha)$ with Diophantine $\alpha$, the theorem of Gottschalk-Hedlund [28] assures that $f$ is a coboundary, meaning $f(x)=F(x+\alpha)-F(x)$, implying that the sum is bounded. Having so proven that the Taylor series has no singularity on $|z|<1$, the partition function $p(n)$ satisfies $\limsup_{n}|p(n)|^{1/n}\geq 1$. The other inequality $\limsup_{n}|p(n)|^{1/n}\leq 1$ implies together with $p(n)\geq 1$ that $\sum_{n=1}^{\infty}p(n)z^{n}\geq\sum_{n=1}^{\infty}z^{n}$ so that the radius of convergence of the former is larger or equal than the radius of convergence of the later, which is $1$. ∎
@@ 5. Psychology
How do human learn, teach and do research? How do they gain intuition about mathematical objects. Psychological aspects of teaching have been looked at in research [20] or education [48]. Hadamard stresses that visual thinking is key for intuition. Norton suggests that a number is defined through our own actions and mentions in the chapter ”What are numbers” cites historical definitions of numbers like Euclid’s notion ”A number is a multitude composed of units.” appearing in book VII. The development of numbers is also linked to language and notation [46]. A more topological approach to numbers are quipu, talking knots [45, 2].
@@
While commutative structures have been challenged since the emergence of quantum mechanics and under the umbrella of non-commutative mathematics [7], the commutativity axiom for the additive structure of the integers has never been put up to debate. But as argued in [40], there are reasons to consider a group generated by two involutions more fundamental than a group by a single generator $1$. The dihedral approach very much is in the philosophy of reflection geometry [3, 24] where reflections are used to generate the non-Abelian Euclidean group $\mathbb{R}^{n}\rtimes O(n)$ consisting of translations and orthogonal transformations. Reflections at affine planes are all involutions. dihedral counting enters geometry if we look at the group generated by the exterior derivative $d$ and its dual $d^{*}$. We have $d^{2}=0$ and $(d^{*})^{2}=0$.
@@
Also notation and implementation is important. Good notation pays off in the long term by allowing to think more effectively. Leibniz’s notations and nomenclature in calculus for example was much persuasive than Newton’s notation. In the case of manipulatives, the analog of notation is how to use material, shape or color and labels implemented in texture. Katherine Stern used colored cubes around a similar time than Cuisenaire. But the colors did not carry intuitive meaning. Other manipulatives had writings on them, making them less elegant.
@@ 6. History
Figure 6. Hardy (played in the movie by Jeremy Irons) explains the asymptotic Hardy-Ramanujan formula for the partition function in the 2015 film: “The man who knew infinity”.
6.5.
@@ 7. Creativity
Partitions appear when defining discrete sub-manifolds of given manifolds [41]. While working on this emerged the picture that the addition of two partitions $n=(n_{1}\dots,n_{k})$, $m=(m_{1},\dots,m_{l})$, in the form of concatenation $n+m=(n_{1},\dots,n_{k},m_{1},\dots,m_{l})$ and the product by foiling out: $n*m=(n_{1}m_{1},n_{2}m_{1},\dots,n_{k}m_{1},\dots,n_{1}m_{k}\dots,n_{k}m_{l})$. For example $(3,4)+(1,1,2)=(3,4,1,1,2)$ and $(3,4)*(1,1,2)=(3,3,6,4,4,8)$. It leads to the idea to associate to a partition $n=(n_{1},\dots,n_{k})$ the $k$-partite graph $K_{n_{1},\dots,n_{k}}$. The partition $5=K_{1,1,1,1,1}$ for example is $K_{5}$ and $6==K_{3,3}$ is the utility graph. The addition is the Zykov join and the multiplication is the Sabidussi multiplication. This ring structure on multi-partite graph is dual to the Shannon ring structure, where addition is disjoint union and multiplication is Shannon multiplication.
@@
The freedom which these teachers had appears more exceptional today. It has become the privilege of private institutions. The times of the “Black Mountain college” [55, 11], where rival methodologies were explored and experimental forms were practiced, are long over. (The college is featured in the 2013 novel ”The longest ride” by Nickolas Sparks, and also made it into a movie in 2015, where the free teaching style at the college is very well visible.)
@@
By the way, the mathematician Max Dehn, who was a student of Hilbert, had taught from 1945-1952 at the Black Mountain college and is buried there [59]. The Wikipedia article about him (which follows [59]) states He enjoyed the forested mountains found in Black Mountain, and would often hold class in the woods, giving lectures during hikes. His lectures frequently drifted off topic on tangents about philosophy, the arts, and nature and their connection to mathematics. He and his wife took part in community meetings and often ate in the dining room. They also regularly had long breakfasts with Buckminster Fuller and his wife. As as source, [50] is given.
@@
The following quotation given in [59] about math education is remarkable: All instruction, including mathematics, should be an end in itself, not, in the first place, means. Therefore in mathematics one should be so simple and elementary as possible so that the structure of mathematical thinking becomes clear. If this is achieved then the student will have less difficulty to see mathematical structures in physical phenomena or even in phenomena of biology and social life. And from [50]: As Dehn was famous for his way of picturing abstract concepts, many of his great mathematical ideas can be intuitively understood through pictures.
Figure 7. This photo credited to the photography critic Nancy Newhall (1908-1974) is titled “Max Dehn at the Black Mountain College” is dated around 1947. The gelatin silver negative had been gifted by Christi Newhall to the Eastman Museum, where one can find a high resolution version. Nancy Newhall visited together with her husband Beaumont Newhall the Black Mountain College for three summers in the years 1946-1949. The Newhalls photographed during that time the campus and its people.
@@
Also in higher education, the administrative cages have been rolled out more and more. We could still develop my calculus courses independently without oversight 10 years ago. Of course, there were course goals and coordination between different courses but the details, the homework or lesson plans were not prescribed. Today, we live in a time, with much more constraints. It appears safer like that but a teaching profession that lacks initiative and originality will be replaced by technology much faster than anticipated.